Determine the probability , Mathematics

Assignment Help:

A medical survey was conducted in order to establish the proportion of the population which was infected along with cancer. The results indicated that 40 percent of the population was suffering from the disease.

A sample of 6 people was later taken and examined for the disease. Determine the probability that the given outcomes were observed

a) Merely one person had the disease

b) Exactly two people had the disease

c) Mostly two people had the disease

d) At least two people had the disease

e) Three or four(4) people had the disease

Solution

P(a persona having cancer) = 40%  = 0.4 = P

P(a person not having cancer) = 60%             = 0.6 = 1 - p = q

a)      P(only one person having cancer)     

= 6C1 (0.4)(0.6)5

=  6!/(5! 1!)(0.4)1(0.6)5                      

= 0.1866

Note that from the formula

nCrprqn-r:           where as: n = sample size = 6

                                    p = 0.4

                                    r = 1 = simply one person having cancer

b)      P(2 people had the disease)

= 6C2 (0.4)2 (0.6)4

6!/(6! 2!)=  (0.4)2 (0.6)5

(6 * 5 * 4!)/(4! * 2 *1)=   (0.4)2 (0.6)5

= 15 × (0.4)2 (0.6)5

= 0.311

c)      P(at most 2) = P(0) + P(1) + P(2) = P(0) or P(1) or P(2)

So we estimate the probability of each and add them up.

P(0) = P(nobody having cancer)

= 6C0 (0.4)0(0.6)6

6!/(0! 6!)=  (0.4)0(0.6)6

= (0.6)6

 = 0.0467

The probabilities of P(1) and P(2) have been worked out in part (a) and

Hence P(at most 2) = 0.0467 + 0.1866 + 0.311 = 0.5443

d)      P(at least 2)

            = P(2) + P(3) + P(4) + P(5) + P(6)

= 1 - [P(0) + P(1)] it is a shorter way of working out the solution as

[P(0) + P(1) + P(2) + P(3) + P(4) + P(5) + P(6) = 1]

= 1 - (0.0467 + 0.1866)

= 0.7667

e)      P(3 or 4 people had the disease)

= P(3) +P(4)

= 6C3(0.4)3(0.6)3  + 6C4(0.4)4(0.6)2

= ( 6!/(3! 3!)) (0.4)3(0.6)3 + (6!/(2! 4!)) (0.4)4(0.6)2

 = {(6 × 5 × 4 × 3!)/ (3 × 2 × 1 × 3!)}(0.4)3(0.6)3 + {(6 × 5 × 4!)/(2 × 1 × 4!)}  (0.4)4(0.6)2

 = 20(0.4)3(0.6)3  + 15(0.4)4(0.6)2 

= (20 × 0.013824) + (15 × 0.009216)

= 0.27648 + 0.13824

 = 0.41472


Related Discussions:- Determine the probability

Regrouping, in regrouping if we have abig number in the end what should i d...

in regrouping if we have abig number in the end what should i do?add an number on top of it,please help

Trigonometry, how to work out consumer arithmetic?

how to work out consumer arithmetic?

Determine the actual viewing area, Computer monitors are calculated by thei...

Computer monitors are calculated by their diagonals. If a monitor is advertised to be 19 in, Determine the actual viewing area, considerthe screen is square? (Round to the nearest

Explain how to distribute simplifying expressions, Explain How to Distribut...

Explain How to Distribute simplifying expressions? The distributive law states that for all numbers a, b, and c, a(b + c)= ab + ac What does this mean in plain language?

Find out solutions to second order differential equations, Find out some so...

Find out some solutions to y′′ - 9 y = 0 Solution  We can find some solutions here simply through inspection. We require functions whose second derivative is 9 times the

Which team should get the ball at the beginning, Why is tossing a coin cons...

Why is tossing a coin considered to be a fair way of deciding which team should get the ball at the beginning of a foot ball match? Ans: equally likely because they are mutual

Proof of limit comparison test - sequences and series, Proof of Limit Compa...

Proof of Limit Comparison Test As 0  Now, as   we know that for large enough n the quotient a n /b n should be close to c and thus there must be a positive integer

Write Your Message!

Captcha
Free Assignment Quote

Assured A++ Grade

Get guaranteed satisfaction & time on delivery in every assignment order you paid with us! We ensure premium quality solution document along with free turntin report!

All rights reserved! Copyrights ©2019-2020 ExpertsMind IT Educational Pvt Ltd